OpenStudy (anonymous):

25x^2 - 60x + 36 = 0

hero (hero):

Hi

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

yeah haha, can you help me?

hero (hero):

Okay

hero (hero):

Hold on a sec

OpenStudy (anonymous):

Thank you (: okay.

hero (hero):

25x^2 - 30x - 30x + 36 = 0 5x(5x-6)-6(5x-6) = 0 (5x-6)(5x-6) = 0

hero (hero):

x = 6/5

OpenStudy (anonymous):

THank you. i was so close.

hero (hero):

Hmmm. Well, if you do it this way, you never get it wrong